### v1

@@ 4.1 Proof of Theorem 1
Consider the graph $G_{1}$ as in Figure 5. Let $B_{1}$ and $B_{2}$ be two isomorphic copies of $G_{1}$. Join $b_{1,21}$ of $B_{1}$ and $b_{2,1}$ of $B_{2}$ by an edge (Figure 6 and 7). We denote this new graph by $G_{2}$ (Figure 6). In this way, we form $G_{k}$ by joining $k$ isomorphic copies of $G_{1}$ : $B_{1},B_{2},\cdots,B_{k}$ (Figure 7). Here $B_{i}$ is joined with $B_{i+1}$ by joining $b_{i,21}$ and $b_{i+1,1}$. We say that $B_{i}$ is the $i$-th block of $G_{k}$. $B_{i}$ is an induced subgraph of $G_{k}$ as given by $B_{i}=G_{k}[\{b_{i,j}:1\leq j\leq 21\}]$. Similarly, for $11\leq i\leq 2k-$ 1, we define $B_{i}\cup B_{i+1}$, induced subgraph of $G_{2k}$, as $B_{i}\cup B_{i+1}=G_{2k}[\{b_{i,j},b_{i+1,j}:1\leq j\leq 21\}]$. We prove Theorem 1 by establishing that $\gamma_{b}(G_{2k})=10k$ and $\operatorname{mp}(G_{2k})=9k$.
@@
Suppose $S_{4}\cap M_{2}\neq\phi$, then $b_{2,7}\notin M_{2}$. Now, $55=|M_{2}|=|M_{2}\cap[V(B_{2})\setminus\{b_{2,7}\}]|=|M_{2}\cap(S^{\prime}\cup S_{4}\cup S_{5}\cup S_{6})|=|(M_{2}\cap S^{\prime})\cup(M_{2}\cap S_{4})\cup(M_{2}\cap S_{5})\cup(M_{2}\cap S_{6})|\leq|M_{2}\cap S^{\prime}|+|M_{2}\cap S_{4}|+|M_{2}\cap S_{5}|+|M_{2}\cap S_{6}|\leq 2+1+1+1=$ 5. Therefore $|S_{t}\cap M_{2}|=1$ for $t=4,5,6$ and $|S^{\prime}\cap M_{2}|=2$. Since $|M_{2}\cap S_{6}|=1$, so there exists $u_{1}\in M_{2}\cap S_{6}$. Then $N_{2}[b_{2,17}]$ contains three vertices $b_{2,14},b_{2,21},u_{1}$ of $M_{2}$, which is not possible. So, this is a contradiction.
Suppose $S_{4}\cap M_{2}=\phi$, then either $b_{2,7}\in M_{2}$ or $b_{2,7}\notin M_{2}$. First consider $b_{2,7}\notin M_{2}$, then $45=|M_{2}|=|M_{2}\cap(S^{\prime}\cup S_{5}\cup S_{6})|=|(M_{2}\cap S^{\prime})\cup(M_{2}\cap S_{5})\cup(M_{2}\cap S_{6})|\leq|M_{2}\cap S^{\prime}|+|M_{2}\cap S_{5}|+|M_{2}\cap S_{6}|\leq 2+1+1=$ 4. So, this is a contradiction. And if $b_{2,7}\in M_{2}$, then $55=|M_{2}|=|M_{2}\cap(S^{\prime}\cup S_{5}\cup S_{6}\cup\{b_{2,7}\})|=|(M_{2}\cap S^{\prime})\cup(M_{2}\cap S_{5})\cup(M_{2}\cap S_{6})\cup(M_{2}\cap\{b_{2,7}\})|\leq|M_{2}\cap S^{\prime}|+|M_{2}\cap S_{5}|+|M_{2}\cap S_{6}|+|M_{2}\cap\{b_{2,7}\}|\leq 2+1+1+1=$ 5. Therefore $|S_{t}\cap M_{2}|=1$ for $t=5,6$ and $|S^{\prime}\cap M_{2}|=2$. Since $|M_{2}\cap S_{6}|=1$, so there exists $u_{2}\in M_{2}\cap S_{6}$. Then $N_{2}[b_{2,17}]$ contains three vertices $b_{2,14},b_{2,21},u_{2}$ of $M_{2}$, which is not possible. So, this is a contradiction. So, $|M_{1}|=5\implies|M_{2}|\leq 4$. $\triangleleft$
@@
Notice that graph $G_{2k}$ has $k$ copies of $G_{2}$. Moreover, we have $\operatorname{mp}(G_{2})=9$. If $\operatorname{mp}(G_{2k})>9k$, then we will use the Pigeonhole principle to show that $\operatorname{mp}(G_{2k})=9k$.
@@
R. C. Brewster and L. Duchesne brewster2013broadcast introduced fractional multipacking in 2013 (also see teshima2014multipackings ). Suppose $G$ is a graph with $V(G)=\{v_{1},v_{2},v_{3},\dots,v_{n}\}$ and $w:V(G)\rightarrow[0,\infty)$ is a function. So, $w(v)$ is a weight on a vertex $v\in V(G)$. Let $w(S)=\sum_{u\in S}w(u)$ where $S\subseteq V(G)$. We say $w$ is a fractional multipacking of $G$, if $w(N_{r}[v])\leq r$ for each vertex $v\in V(G)$ and for every integer $r\geq 1$. The fractional multipacking number of $G$ is the value $\max_{w}w(V(G))$ where $w$ is any fractional multipacking and it is denoted by $\operatorname{mp}_{f}(G)$. A maximum fractional multipacking is a fractional multipacking $w$ of a graph $G$ such that $w(V(G))=\operatorname{mp}_{f}(G)$. If $w$ is a fractional multipacking, we define a vector $y$ with the entries $y_{j}=w(v_{j})$. So,
@@
Define a broadcast ${f}$ on $G_{k}$ as ${f}(b_{i,j})=\begin{cases}2&\text{if }1\leq i\leq k\text{ and }j=6,17\\ 1&\text{if }1\leq i\leq k\text{ and }j=12\\ 0&\text{otherwise }\end{cases}$.
Here ${f}$ is an efficient dominating broadcast and $\sum_{v\in V(G_{k})}{f}(v)=5k$. So, $\gamma_{b}(G_{k})\leq 5k$, $\forall k\in\mathbb{N}$. So, by the strong duality theorem and Lemma 6, $5k\leq\operatorname{mp}_{f}(G_{k})=\gamma_{b,f}(G_{k})\leq\gamma_{b}(G_{k})\leq 5k$. Therefore, $\operatorname{mp}_{f}(G_{k})=\gamma_{b}(G_{k})=5k$. ∎
@@ 5 Hyperbolic graphs
Chepoi et al. chepoi2008diameters established a relation between the length of the radius and diameter of a $\delta$-hyperbolic graph, that we state in the following theorem:
Theorem 5 (chepoi2008diameters ).
@@
Let $G$ be a $\delta$-hyperbolic graph with radius $r$ and diameter $d$. From Theorem 2, $\big{\lceil}{\frac{d+1}{3}\big{\rceil}}\leq\operatorname{mp}(G)$ which implies that $d\leq 3\operatorname{mp}(G)-1$. Moreover, from Theorem 3 and Theorem 5, $\gamma_{b}(G)\leq r\leq\big{\lfloor}{\frac{d+4\delta+1}{2}\big{\rfloor}}\leq\big{\lfloor}{\frac{(3\operatorname{mp}(G)-1)+4\delta+1}{2}\big{\rfloor}}$ $=\big{\lfloor}{\frac{3}{2}\operatorname{mp}(G)+2\delta\big{\rfloor}}$. Therefore, $\gamma_{b}(G)\leq\big{\lfloor}{\frac{3}{2}\operatorname{mp}(G)+2\delta\big{\rfloor}}$. ∎
@@ 6 Conclusion
We have shown that the bound $\gamma_{b}(G)\leq 2\operatorname{mp}(G)+3$ for general graphs $G$ can be improved to $\gamma_{b}(G)\leq\big{\lceil}{\frac{3}{2}\operatorname{mp}(G)\big{\rceil}}$ for connected chordal graphs. It is known that for strongly chordal graphs, $\gamma_{b}(G)=\operatorname{mp}(G)$, we have shown that this is not the case for connected chordal graphs. Even more, $\gamma_{b}(G)-\operatorname{mp}(G)$ can be arbitrarily large for connected chordal graphs, as we have constructed infinitely many connected chordal graphs $G$ where $\gamma_{b}(G)/\operatorname{mp}(G)=10/9$ and $\operatorname{mp}(G)$ is arbitrarily large. Moreover, we have shown that $\gamma_{b}(G)\leq\big{\lfloor}{\frac{3}{2}\operatorname{mp}(G)+2\delta\big{\rfloor}}$ holds for the $\delta$-hyperbolic graphs. Note that connected split graphs have radius at most 2, so their broadcast number and their multipacking number are both at most 2. The graph from Figure 2 is a split graph, showing that the multipacking number of a split graph can indeed be different from its broadcast domination number.
It remains an interesting open problem to determine the best possible value of $\lim_{\operatorname{mp}(G)\to\infty}\sup\bigg{\{}\displaystyle\frac{\gamma_{b}(G)}{\operatorname{mp}(G)}\bigg{\}}$ for general connected graphs and for connected chordal graphs. This problem could also be studied for other interesting graph classes such as the one in Figure 1, or others.

### v2

@@ 4.1 Proof of Theorem 1
Consider the graph $G_{1}$ as in Figure 5. Let $B_{1}$ and $B_{2}$ be two isomorphic copies of $G_{1}$. The vertices of $B_{1}$ and $B_{2}$ are labeled as the vertices of $B_{i}$ and $B_{i+1}$ (for $i=1$) in Fig 7. Join $b_{1,21}$ of $B_{1}$ and $b_{2,1}$ of $B_{2}$ by an edge (Figure 6 and 7). We denote this new graph by $G_{2}$ (Figure 6). In this way, we form $G_{k}$ by joining $k$ isomorphic copies of $G_{1}$ : $B_{1},B_{2},\cdots,B_{k}$ (Figure 7). Here $B_{i}$ is joined with $B_{i+1}$ by joining $b_{i,21}$ and $b_{i+1,1}$. We say that $B_{i}$ is the $i$-th block of $G_{k}$. $B_{i}$ is an induced subgraph of $G_{k}$ as given by $B_{i}=G_{k}[\{b_{i,j}:1\leq j\leq 21\}]$. Similarly, for $1\leq i\leq 2k-1$, we define $B_{i}\cup B_{i+1}$, induced subgraph of $G_{2k}$, as $B_{i}\cup B_{i+1}=G_{2k}[\{b_{i,j},b_{i+1,j}:1\leq j\leq 21\}]$. We prove Theorem 1 by establishing that $\gamma_{b}(G_{2k})=10k$ and $\operatorname{mp}(G_{2k})=9k$.
@@
Suppose $S_{4}\cap M_{2}\neq\phi$, then $b_{2,7}\notin M_{2}$. Now, $5=|M_{2}|=|M_{2}\cap[V(B_{2})\setminus\{b_{2,7}\}]|=|M_{2}\cap(S^{\prime}\cup S_{4}\cup S_{5}\cup S_{6})|=|(M_{2}\cap S^{\prime})\cup(M_{2}\cap S_{4})\cup(M_{2}\cap S_{5})\cup(M_{2}\cap S_{6})|\leq|M_{2}\cap S^{\prime}|+|M_{2}\cap S_{4}|+|M_{2}\cap S_{5}|+|M_{2}\cap S_{6}|\leq 2+1+1+1=5$. Therefore $|S_{t}\cap M_{2}|=1$ for $t=4,5,6$ and $|S^{\prime}\cap M_{2}|=2$. Since $|M_{2}\cap S_{6}|=1$, so there exists $u_{1}\in M_{2}\cap S_{6}$. Then $N_{2}[b_{2,17}]$ contains three vertices $b_{2,14},b_{2,21},u_{1}$ of $M_{2}$, which is not possible. So, this is a contradiction.
Suppose $S_{4}\cap M_{2}=\phi$, then either $b_{2,7}\in M_{2}$ or $b_{2,7}\notin M_{2}$. First consider $b_{2,7}\notin M_{2}$, then $5=|M_{2}|=|M_{2}\cap(S^{\prime}\cup S_{5}\cup S_{6})|=|(M_{2}\cap S^{\prime})\cup(M_{2}\cap S_{5})\cup(M_{2}\cap S_{6})|\leq|M_{2}\cap S^{\prime}|+|M_{2}\cap S_{5}|+|M_{2}\cap S_{6}|\leq 2+1+1=4$. So, this is a contradiction. And if $b_{2,7}\in M_{2}$, then $5=|M_{2}|=|M_{2}\cap(S^{\prime}\cup S_{5}\cup S_{6}\cup\{b_{2,7}\})|=|(M_{2}\cap S^{\prime})\cup(M_{2}\cap S_{5})\cup(M_{2}\cap S_{6})\cup(M_{2}\cap\{b_{2,7}\})|\leq|M_{2}\cap S^{\prime}|+|M_{2}\cap S_{5}|+|M_{2}\cap S_{6}|+|M_{2}\cap\{b_{2,7}\}|\leq 2+1+1+1=5$. Therefore $|S_{t}\cap M_{2}|=1$ for $t=5,6$ and $|S^{\prime}\cap M_{2}|=2$. Since $|M_{2}\cap S_{6}|=1$, so there exists $u_{2}\in M_{2}\cap S_{6}$. Then $N_{2}[b_{2,17}]$ contains three vertices $b_{2,14},b_{2,21},u_{2}$ of $M_{2}$, which is not possible. So, this is a contradiction. So, $|M_{1}|=5\implies|M_{2}|\leq 4$. $\triangleleft$
@@
Notice that graph $G_{2k}$ has $k$ copies of $G_{2}$. Moreover, we have $\operatorname{mp}(G_{2})=9$. Using the Pigeonhole principle, we show that $\operatorname{mp}(G_{2k})=9k$.
@@
R. C. Brewster and L. Duchesne [4] introduced fractional multipacking in 2013 (also see [29]). Suppose $G$ is a graph with $V(G)=\{v_{1},v_{2},v_{3},\dots,v_{n}\}$ and $w:V(G)\rightarrow[0,\infty)$ is a function. So, $w(v)$ is a weight on a vertex $v\in V(G)$. Let $w(S)=\sum_{u\in S}w(u)$ where $S\subseteq V(G)$. We say $w$ is a fractional multipacking of $G$, if $w(N_{r}[v])\leq r$ for each vertex $v\in V(G)$ and for every integer $r\geq 1$. The fractional multipacking number of $G$ is the value $\max_{w}w(V(G))$ where $w$ is any fractional multipacking and it is denoted by $\operatorname{mp}_{f}(G)$. A maximum fractional multipacking is a fractional multipacking $w$ of a graph $G$ such that $w(V(G))=\operatorname{mp}_{f}(G)$. If $w$ is a fractional multipacking, we define a vector $y$ with the entries $y_{j}=w(v_{j})$. So,
@@
Define a broadcast f{f} on $G_{k}$ as ${f}(b_{i,j})=\begin{cases}2&\text{if }1\leq i\leq k\text{ and }j=6,17\\ 1&\text{if }1\leq i\leq k\text{ and }j=12\\ 0&\text{otherwise }\end{cases}$.
Here f{f} is an efficient dominating broadcast and $\sum_{v\in V(G_{k})}{f}(v)=5k$. So, $\gamma_{b}(G_{k})\leq 5k$, $\forall k\in\mathbb{N}$. So, by the strong duality theorem and Lemma 6, $5k\leq\operatorname{mp}_{f}(G_{k})=\gamma_{b,f}(G_{k})\leq\gamma_{b}(G_{k})\leq 5k$. Therefore, $\operatorname{mp}_{f}(G_{k})=\gamma_{b}(G_{k})=5k$. ∎
@@ 5 Hyperbolic graphs
Chepoi et al. [9] established a relation between the length of the radius and diameter of a $\delta$-hyperbolic graph, that we state in the following theorem:
Theorem 5 ([9]).
@@
Let $G$ be a $\delta$-hyperbolic graph with radius $r$ and diameter $d$. From Theorem 2, $\big\lceil{\frac{d+1}{3}\big\rceil}\leq\operatorname{mp}(G)$ which implies that $d\leq 3\operatorname{mp}(G)-1$. Moreover, from Theorem 3 and Theorem 5, $\gamma_{b}(G)\leq r\leq\big\lfloor{\frac{d+4\delta+1}{2}\big\rfloor}\leq\big\lfloor{\frac{(3\operatorname{mp}(G)-1)+4\delta+1}{2}\big\rfloor}$ $\big\lfloor{\frac{3}{2}\operatorname{mp}(G)+2\delta\big\rfloor}$. Therefore, $\gamma_{b}(G)\leq\big\lfloor{\frac{3}{2}\operatorname{mp}(G)+2\delta\big\rfloor}$. ∎
@@ 6 Conclusion
We have shown that the bound $\gamma_{b}(G)\leq 2\operatorname{mp}(G)+3$ for general graphs $G$ can be improved to $\gamma_{b}(G)\leq\big\lceil{\frac{3}{2}\operatorname{mp}(G)\big\rceil}$ for connected chordal graphs. It is known that for strongly chordal graphs, $\gamma_{b}(G)=\operatorname{mp}(G)$, we have shown that this is not the case for connected chordal graphs. Even more, $\gamma_{b}(G)-\operatorname{mp}(G)$ can be arbitrarily large for connected chordal graphs, as we have constructed infinitely many connected chordal graphs $G$ where $\gamma_{b}(G)/\operatorname{mp}(G)=10/9$ and $\operatorname{mp}(G)$ is arbitrarily large. Moreover, we have shown that $\gamma_{b}(G)\leq\big\lfloor{\frac{3}{2}\operatorname{mp}(G)+2\delta\big\rfloor}$ holds for all $\delta$-hyperbolic graphs. Note that connected split graphs have radius at most 2, so their broadcast number and their multipacking number are both at most 2. The graph from Figure 2 is a split graph, showing that the multipacking number of a split graph can indeed be different from its broadcast domination number.
We naturally ask for the best possible ratio on chordal graphs, which our work leaves open. This problem could also be studied for other interesting graph classes, such as the one in Figure 1, or others.
For general connected graphs, although the optimal ratio is now known asymptotically since ${\lim_{\operatorname{mp}(G)\to\infty}\sup\left\{\frac{\gamma_{b}(G)}{\operatorname{mp}(G)}\right\}=2}$, with a lower bound coming from hypercubes [26], it remains an interesting open problem to tighten this ratio further. One direction is to show that $\gamma_{b}(G)\leq 2\operatorname{mp}(G)$ holds. For the lower bound, it would be nice to have constructions of connected graphs $G$ with arbitrarily large values of $\operatorname{mp}(G)$ that attain the ratio of 2. In fact, the exact values of $\operatorname{mp}(H_{d})$ are still open [26] and, perhaps, they come close to a ratio of 2?
Declarations
Funding: This research was financed by the IFCAM project “Applications of graph homomorphisms” (MA/IFCAM/18/39). Florent Foucaud was financed by the French government IDEX-ISITE initiative 16-IDEX-0001 (CAP 20-25), the International Research Center ”Innovation Transportation and Production Systems” of the I-SITE CAP 20-25, the ANR project GRALMECO (ANR-21-CE48-0004), and the CNRS IRL RELAX during the preparation of this manuscript.